OpenStudy (anonymous):

Write each linear equation. (Use point point-slope form.) slope = 1/2 Passes through (2,3) What do I do?

OpenStudy (anonymous):

Well the equation for point slope is y-y1=m(x-x1) Using (x,y) and m as the slope. Your slope is 1/2 Your x is 2, and your y is 3. (x,y)(2,3) So y-3=1/2(x-2) Distribute the 1/2 y-3=1/2x-1 Add the 3 to both sides. y=1/2x+2

OpenStudy (anonymous):

Thanks!